OpenStudy (anonymous):

a broken clover has 2 leaves, a normal clover has 3 leaves, and a lucky clover has 4 leaves. in a bunch of these clovers, there are twice as many normal clovers as broken clovers and 5 times as many normal clovers as lucky clovers. the bunch has a total of N leaves. N is greater than 200. what is the least value of N?

pooja195 (pooja195):

If X = number of broken clovers 2X + 2 (3X) + 2/5 (4X) = N 10X + 30X + 8X = 5N 48X = 5N X = 5N / 48 Possible N = 5 x 48 = 240 X = 5 x 5 = 25 Answer X = 25

OpenStudy (anonymous):

the question is to find of the value of N not X, so the answer is N=240?

OpenStudy (anonymous):

but thank you for your help

pooja195 (pooja195):

yr welcome but do u get this

OpenStudy (anonymous):

just stil confusing

pooja195 (pooja195):

srry but not at my skill

OpenStudy (anonymous):

ok i will ask to other members @jim_thompson5910

OpenStudy (anonymous):

@kropot72

jimthompson5910 (jim_thompson5910):

pooja195 is correct You want to make x = 5N/48 equal to some integer 200/48 = 4 remainder 8 so 192/48 = 4 remainder 0 (subtract 8 from 200 to get 192) Now add 48 to 192 to get 192+48 = 240 This is the smallest value of N possible to make 5N/48 So that's why the answer is 240

jimthompson5910 (jim_thompson5910):

*to make 5N/48 an integer

OpenStudy (anonymous):

thank you sir!

jimthompson5910 (jim_thompson5910):

np